OpenStudy (anonymous):

Number of ways the word 'SUCCESS' can be arranged, such that no two S's and C's are together.

OpenStudy (anonymous):

@Callisto @lgbasallote @amistre64 @shubhamsrg @Shane_B

OpenStudy (anonymous):

@waterineyes ....Can u Solve it out!

OpenStudy (anonymous):

96

OpenStudy (anonymous):

\[\binom{5}3\frac{4!}{2!}-\binom433!\]

OpenStudy (anonymous):

Did nt understand....

OpenStudy (anonymous):

or The total number of permutation of letters (T)= 7!2!3! With two cc together (A)= 6!2! With three ss together (B)= 6!2!−5!2! With both ss and cc together (C)= 5!−4! Answer = T - A - B + C = 96

OpenStudy (anonymous):

@Derp_Derpington Plzz mention the site from which u copied this..... http://math.stackexchange.com/questions/162394/arrangement-of-the-word-success

OpenStudy (anonymous):

It is against the COD if u copy paste the answers

OpenStudy (anonymous):

oh yeah sorry forgot to

OpenStudy (anonymous):

thats ok...

OpenStudy (anonymous):

@waterineyes plzz expl this.....i did nt understand

OpenStudy (anonymous):

That will be 7!/2! 3! @Derp_Derpington right??

OpenStudy (anonymous):

This site is not working properly for me @Yahoo!

OpenStudy (shubhamsrg):

is ans 96 only ?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73

OpenStudy (anonymous):

@amistre64 @amistre64 @amistre64 @amistre64

OpenStudy (anonymous):

@waterineyes yes it is!

OpenStudy (amistre64):

im thinking a tree diagram would be rather long and involved eh

OpenStudy (amistre64):

success; 7 tiers, 4 letters 3s1u2c1e s1 u c1 e c1 e s2 u s2 e u s2 c1 es2 c1s2 ec1 c2s2e c2ue uc2s2 s2c2 uc2 s2u it might go simpler on paper tho

OpenStudy (anonymous):

wat is this

OpenStudy (amistre64):

a tree diagram .... which was the idea i proposed :/

OpenStudy (amistre64):

working it out on paper might go easier than trying to type it up readably on the computer screen tho